OpenStudy (anonymous):

I need help figuring out how to do this compound inequality. 4 ≤ 3x + 1 ≤ 9

pooja195 (pooja195):

Our goal is to get x alone. What would be the first step?

OpenStudy (anonymous):

since it says to add 1 do you subtract the one from the 9?

pooja195 (pooja195):

Correct we subtract 1 HOWEVER what we do to one number we do to all Like this: 

pooja195 (pooja195):

4-1=?

OpenStudy (anonymous):

3

pooja195 (pooja195):

1-1=?

OpenStudy (anonymous):

0

pooja195 (pooja195):

Good so it cancels 9-1=?

OpenStudy (anonymous):

8

pooja195 (pooja195):

Good now we have 3 ≤ 3x ≤ 8 Whats next?

pooja195 (pooja195):

Any ideas?

OpenStudy (anonymous):

that's where i'm stuck

OpenStudy (anonymous):

wait do i add the 3 and 8 together?

pooja195 (pooja195):

Well rememeber that we want to get x alone how would we do that? What would we have to do with the 3 hint it would be the opposite of multiplcation.

pooja195 (pooja195):

No we would not add

OpenStudy (anonymous):

divide

pooja195 (pooja195):

Perfect  What would we be left with?